# Magic phase transition and non-local complexity in generalized $W$ State

### v1

@@ Appendix A Magic phase transition and non-local complexity in generalized W State – Supplementary Meterials
$\lambda_{1,\ldots,4}=\frac{1}{4L}\Big{(}L+2\gamma\cos(p\chi)\pm\sqrt{(L-2a)^{2}+4L(1+\gamma\cos(p\chi))-4\sin^{2}(p\chi)}\Big{)}\,,$ (16)
@@
$S_{2}(a,p)=-\log_{2}\left[\frac{L(2+L)-2a(L-a)+2\cos(p\chi)}{2L^{2}}\right]$ (17)
@@
A.2 Analytic derivations of Stabilizer Rényi entropy for a $W_{P}$ state
The main result shown in the main part of the letter is based on the behavior of the SRE on the family of $W_{P}$ states, which is a family that generalizes the well-known $W$ state. To evaluate the SRE, let us start by recalling its expression for a generic pure state |$\ket{\psi}$. It reads
@@ A.2 Analytic derivations of Stabilizer Rényi entropy for a $W_{P}$ state
The second exception is represented by the Pauli strings with only two operators in the set $\{\sigma^{y},\,\sigma^{z}\}$. Within this hypothesis, the Pauli string can be written as $\mathcal{P}^{\prime\prime}=\bigotimes_{k=1,k\neq i,j}^{L}\sigma_{k}^{\alpha}\otimes(\sigma^{\beta}_{j}\sigma^{\gamma}_{k})$ where $\alpha=0,x$ while $\beta,\,\gamma=y,z$. This contribution comes from the fact that such strings are able to shift the $\ket{+}$ from the site $j$ to the site $k$ and vice versa. When $\beta=\gamma$ both these two terms have the same sign, so giving a contribution proportional to $\cos{[(j-k)p]}/2L$. On the contrary, when $\beta\neq\gamma$ they show opposite signs, so contributing proportional to $\sin[(j-k)p]/2L$ Naming $r=j-k$ and summing over all possible Pauli strings we have
$\sum_{\mathcal{P}^{\prime\prime}}O(\mathcal{P}^{\prime\prime})=L\sum_{l=0}^{L-2}\sum_{r=1}^{L-1}\frac{(L-2)!}{l!(L-2-l)!}\left[\left(\frac{2\cos(pr)}{L}\right)^{4}+\left(\frac{2\sin(pr)}{L}\right)^{4}\right].$ (22)
Putting the two non-vanishing contributions together in the definition of the SRE of order 2 we recover, after some steps, the following expression

### v2

@@ Abstract
$\lambda_{1,\ldots,4}$ $=$ $\frac{1}{4L}\Big{(}L+2\gamma\cos(pa)\pm\Big{.}$
±.$\pm\Big{.}\sqrt{(L-2a)^{2}+4L(1+\gamma\cos(pa))-4\sin^{2}(pa)}\Big{)},$
@@
$\!\!\!\!\!\!S_{2}(a,p)\!=\!-\log_{2}\!\left[\frac{L(2+L)\!-\!2a(L-a)\!+\!2\cos(pa)}{2L^{2}}\right],$ (23)
@@
Appendix B Analytic derivations of Stabilizer Rényi entropy for a $W_{P}$ state
The main result shown in the main part of the paper is based on the behavior of the SRE on the family of $W_{P}$ states, which is a family that generalizes the well-known $W$ state. To evaluate the SRE, let us start by recalling its expression for a generic pure state |$\ket{\psi}$. It reads
@@ Appendix B Analytic derivations of Stabilizer Rényi entropy for a $W_{P}$ state
The second exception is represented by the Pauli strings with only two operators in the set $\{\sigma^{y},\,\sigma^{z}\}$. Within this hypothesis, the Pauli string can be written as $\mathcal{P}^{\prime\prime}=\bigotimes_{k=1,k\neq i,j}^{L}\sigma_{k}^{\alpha}\otimes(\sigma^{\beta}_{j}\sigma^{\gamma}_{k})$ where $\alpha=0,x$ while $\beta,\,\gamma=y,z$. This contribution comes from the fact that such strings are able to shift the $\ket{+}$ from the site $j$ to the site $k$ and vice versa. When $\beta=\gamma$ both these two terms have the same sign, so giving a contribution proportional to $\cos{[(j-k)p]}/2L$. On the contrary, when $\beta\neq\gamma$ they show opposite signs, so contributing proportional to $\sin[(j-k)p]/2L$. Naming $r=j-k$ and summing over all possible Pauli strings we have
$\sum_{\mathcal{P}^{\prime\prime}}O(\mathcal{P}^{\prime\prime})$ $=$ $L\sum_{l=0}^{L-2}\sum_{r=1}^{L-1}\frac{(L-2)!}{l!(L-2-l)!}$
$\times$ $\left[\left(\frac{2\cos(pr)}{L}\right)^{4}+\left(\frac{2\sin(pr)}{L}\right)^{4}\right].$
Putting the two non-vanishing contributions together in the definition of the SRE of order 2, we recover, after some steps, the following expression
@@
Appendix C Incompatible symmetries
Figure 4: Left Panel: Dependence on the position of the entanglement properties for the states |$\ket{\phi(p,\theta)}$ in (C). The entanglement is evaluated considering a partition (A|B) where $A$ is made by $(L-1)/2$ contiguous spins of which the first is $k^{*}$, $p=2\pi/L$ and $\theta=(L-1)/4$. Left Panel: Dependence of the amplitude of the oscillations of the entanglement properties on the dimension of the systems. The difference between the maximum and the minimum of $S(A,k^{*})$ is plotted as a function of $1/L$ to highlight the fact that in the thermodynamic limit it vanishes.
In our paper, we considered the breaking of the mirror symmetry and observed how the violation of this symmetry, associated with the emergence of finite-momentum ground states, is evidenced by a finite jump in the value of the SRE. At the same time, we also observed how entanglement does not show any response to this phenomenon and, therefore, is not a quantity usable to detect the breaking of this symmetry and the quantum phase transition it entails. However, the phenomenology connected to this phase transition is very complex and has other aspects to highlight.
In fact, among others, the XYZ chain has two symmetries that are usually considered as trivial, namely the translational symmetry and the mirror symmetry that exchanges sites at the same distance from a reference point. In general, these two symmetries cannot be diagonalized simultaneously, and only states with zero or $\pi$ momentum can be eigenstates of both. However, the $\pi$ momentum is incompatible with the constraint of an odd number of spins.
Crossing the phase transition at $h=h^{*}$ we have a sudden change in ground state degeneracy connected with the appearance of ground states with finite and opposite momenta. In a way, the peculiarity of this phase transition, which renders it different from the usual ones, is that, after crossing it, it is not possible to select out of the degenerate ground state manifold a state that respects all the symmetries of the Hamiltonian.
Figure 5: Dependence on the position of the entanglement properties for the ground states that preserve the mirror symmetry for different system lengths. The ground states considered in this plot are obtained by the DMRG algorithm setting $J_{x}=1$ and $J_{y}=0.33$. With these parameters, the ground states analyzed are the symmetric linear combination of the two ground states with momentum equal to $p=2\pi/L$. The entanglement is obtained considering a partition (A|B) where $A$ is made by $(L-1)/2$ contiguous spins of which the first is $k^{*}$.
In the main part of this paper, we consider eigenstates of the momentum operator that thus break mirror symmetry: let us now look at eigenstates of the mirror symmetry, constructed as the symmetric superposition of states with opposite momenta:
|$\ket{\phi(p,\theta)}$ $=$ $\frac{1}{\sqrt{2}}(e^{-\imath\theta}\ket{\omega(p)}+e^{\imath\theta}\ket{\omega(-p)})$
$=$ $\frac{1}{\sqrt{L}}\sum_{k=1}^{L}\cos(pk-\theta)(\ket{k}+\ket{k^{\prime}})$
It is easy to see that |$\ket{\phi(p,\theta)}$ does not violate the mirror symmetry with respect to the site $\tilde{k}=\theta/p$, but it is not an eigenstate of the spatial translation invariance symmetry of the Hamiltonian, as can be easily seen from the explicit dependence on the modulus $\cos(pk-\theta)$ from the site index $k$.
As we can appreciate from the plot in fig. 4, the entanglement properties of the state |$\ket{\phi(p,\theta)}$ depend on the value $p$ but this dependence vanishes while the dimension of the system diverges. Indeed, in the first panel of fig. 4 we illustrate the dependence of the Von Neumann entropy for a convex bipartition $(A|B)$, where $A$ consists of $(L-1)/2$ contiguous spins, on $k^{*}$ that represents the first spin of the chain in $A$. The analysis is performed for chains of varying lengths, assuming $p=2\pi/L$ and $\theta=(L-1)/4$. As the size of the system increases, the amplitude of the oscillations in the entanglement value decreases, vanishing in the thermodynamic limit, as can be appreciated in the second panel of the same figure. A numerical analysis of the data shows how, for $p=2\pi/L$, in the limit of large $L$, this amplitude can be approximated by $const/L$ where $const\simeq 4.17$. This picture does not change significantly moving away from this perturbative regime: a similar behavior can be be seen in fig. 5 for the entanglement immediately below the transition point for the ground state of the spin chain obtained from DMRG.
Finally, we should mention that the XYZ chain also possesses a $\mathbb{Z}_{2}$ symmetry, related to the spin flip symmetry, and represented by the parity of the magnetization along the $z$-axis. For $h>h^{*}$ the ground state has a fixed parity and is separated from a state with opposite parity by a gap closing polynomially with the system size [26] (this should be contrasted with the case without topological frustration, where the gap closes exponentially [53]). Below the phase transition ($h<h^{*}$), for finite chains the lowest energy states can have different parities, with frequent level crossings between them as the Hamiltonian parameters are changed (the first of which always being at $h=h^{*}$) and for different chain lengths $L$. Overall, the gap between opposite parities’ lowest energy states closes exponentially [29], indicating that in the thermodynamic limit, the ground state manifold doubles its degeneracy. Hence, in addition to ground states breaking the mirror and/or translation symmetry, one could also select a state with indefinite parity to study the phase transition properties. Such a choice introduces further complications and subtleties in the analysis, since the momenta of states with different parities are shifted by $\pi$ and thus several inequivalent scenarios arise, depending on the ground state choice. We believe that these subtleties are peculiar to the model under consideration and will be the subject of a separate work with detailed considerations on different finite-size scenarios, while the SRE discontinuity we reported in the main part of this paper is a general phenomenon that can help in detecting novel types of phase transitions.